OpenStudy (anonymous):

andrew thinks he has a good thing going. he invests 8000 dollars with the firm dewey, cheatham, and howe, purchasing a certificate of deposit that after 10 years wil be worth 10400 dollars. the trouble is andrew did not consult his investment advisor before putting his own money down, and this sneaky company has locked him into an investment that pay simple interest not compound on his principal. what percent is the companys yearly interest rate?

ganeshie8 (ganeshie8):

which part you want help with.. ?

OpenStudy (anonymous):

i neion or something @ganeshie8

ganeshie8 (ganeshie8):

hmm... dint get u..

OpenStudy (anonymous):

oh sorry wow my comp bud out i need an equation or something because i have no idea how to get to the answer @ganeshie8

jimthompson5910 (jim_thompson5910):

The interest is 10400 - 8000 = 2400 dollars The principle is 8000 dollars. The rate is unknown The time is 10 years So I = 2400 P = 8000 r = x t = 10

jimthompson5910 (jim_thompson5910):

Now plug them into the formula I = Prt to get 2400 = 8000*x*10 then solve for x to find your answer.

OpenStudy (anonymous):

sooo youre left with 33 after you divid both sides? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Divide both sides by what?

OpenStudy (anonymous):

to find x dont you have to divide by 2400

jimthompson5910 (jim_thompson5910):

No, you divide both sides by 80000 2400 = 8000*x*10 2400 = 8000*10*x 2400 = 80000*x 2400/80000 = x

jimthompson5910 (jim_thompson5910):

I'll let you finish it

OpenStudy (anonymous):

so you get .03 how do i translate that to a percent

jimthompson5910 (jim_thompson5910):

To convert 0.03 to a percent, move the decimal over to the right 2 spots So 0.03 converts to 3%

OpenStudy (anonymous):

ooohh alright